OpenStudy (anonymous):

3x-4y=10 2x-3y=7 use the elimination system

OpenStudy (anonymous):

3x-4y=10 2x-3y=7 equiv a 9x-12y=30 8x-12y=28 9x-12y-(8x-12y)=x=2

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 3x-4y=10&{\color{red}{ \times -3}}\implies &-9x+12y=-30\\ 2x-3y=7&{\color{red}{ \times 4}}\implies &\quad 8x-12y=28\\ \hline\\ &&-x\qquad \qquad =-18 \end{array}\)

OpenStudy (texaschic101):

3x - 4y = 10 --->(3) 2x - 3y = 7 --->(-4) ---------------- 9x - 12y = 30 (result of multiplying by 3) -8x + 12y = - 28 (result of multiplying by -4) ----------------add x = 2 now sub 2 in for x in either equation to find y 3x - 4y = 10 3(2) - 4y = 10 6 - 4y = 10 -4y = 10 - 6 -4y = 4 y = -1 check.. 2x - 3y = 7 2(2) - 3(-1) = 7 4 + 3 = 7 7 = 7 (correct) solution (2,-1) any questions ?

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 3x-4y=10&{\color{red}{ \times -3}}\implies &-9x+12y=-30\\ 2x-3y=7&{\color{red}{ \times 4}}\implies &\quad 8x-12y=28\\ \hline\\ &&-x\qquad \qquad =-2 \end{array}\) rather... anyhow so you'd substitute that value on either equation to find "y" as shown already